A 6kg clay ball is directly against a perpendicular brick wall at a velocity of 22m/s and shatters in to pieces, which all fly backward. the wall exerts a force on the ball of 2640 N for 0.1s. one piece of mass 2 kg travel backward at a velocity of 10m/s and an angle 32 above the horizontal. A second piece of mass 1 kg travels at a velocity of 8m/s and an angle of 28 below the horizontal. what is the velocity of the third piece?
1
Expert's answer
2021-09-14T14:15:59-0400

"F_x=2640\ (N)\ \ \ and \ \ \ F_y=0"


So,


"0=2\cdot10\cdot\sin32°-1\cdot8\cdot\sin28°+3\cdot v_y\to v_y=-6.84\ (m/s)"


"2640\cdot0.1=(-2\cdot10\cdot\cos32°-1\cdot8\cdot\cos28°-3\cdot v_x)-6\cdot22\to v_x=140\ (m/s)"


"v=\sqrt{140^2+6.84^2}=140.2\ (m/s)"


"\tan\alpha=6.84/140=0.488\to\alpha=2.79°" (below the horizontal)